# Deep Backtracking Counterfactuals for Causally Compliant Explanations

### v3

@@ A.2 Formal Derivation of DeepBC
We derive 5 and 3 from the three-step-procedure of backtracking counterfactuals (see § A.1) as follows:
@@
    We observe that 16 corresponds to performing stochastic DeepBC 4.
@@
    We conclude by noting that $\text{arg}\;\underset{\mathbf{u}^{\prime}}{\text{max}}\;\text{log}\;p(\mathbf{u}^{\prime}\;|\;\mathbf{F}^{-1}(\mathbf{x}),\mathbf{x}^{*}_{S})$, composed with $\mathbf{F}$, is equivalent to 5.
A.3 Derivation of 5
As a result of the linearization of $\mathbf{F}$, $\mathcal{L}(\mathbf{u}^{\prime};\mathbf{u},\mathbf{x}^{*}_{S})$ in 9 simplifies to
@@ B.1 Technical Details and Comments for the DeepBC Optimization Algorithm
In practice, we implement 12 as follows
@@
where $\dagger$ denotes Moore-Penrose pseudoinverse. We employ 18 rather than 12 for the reason of numerical stability. The main computational bottleneck in Alg. 1 is the computation of the pseudoinverse $(\lambda^{-1}\mathbf{W}+\mathbf{J}_{S}^{\top}\mathbf{J}_{S})^{\dagger}$ in 18, which comes at a cost of $\mathcal{O}(\#\text{it}\cdot\text{dim}(\mathbf{u})^{3})$, compared to $\mathcal{O}(\#\text{it}\cdot\text{dim}(\mathbf{u}))$ for gradient descent. We note, however, that the dimensionality of the latent space is typically not very large in our experiments. The maximum dimension is $516$ for CelebA, due to $4$ attributes and $512$-dimensional latent space of the VAE. We also stress that $\mathbf{J}_{S}$ is sparse (many 00 entries) when $S$ covers attribute variables, because the $512$-dimensional latent vector is not upstream of any attribute. We do not run experiments where many variables are upstream of the antecedent variable and stress that this may affect the performance of Alg. 1.
@@
Adam’s convergence highly depends on the choice of learning rate. We suspect that this is due to the poorly conditioned Hessian that comes from choosing large $\lambda$. However, large $\lambda$ is required in order to (at least approximately) fulfill the constraint in 5 (see top row in the table of Fig. 9).
Figure 9: The figures show the sum of penalty losses 9 over all points in Fig. 4 (b) for optimizing it for $200$ iterations on a $\text{log}_{10}$ scale. Comparison of the Adam optimizer with various learning rates in comparison to constraint linearization (Alg. 1) for different choices of penalty parameter $\lambda$. The table shows the sum of constraint errors $\left\lVert\mathbf{F}_{S}(\mathbf{u}^{\prime})-\mathbf{x}^{*}_{S}\right\rVert_{2}^{2}$ after $200$ iterations (How well the constraints are fulfilled).
@@ Appendix C Ground Truth Structural Equations in Morpho-MNIST
C.1 Metric Implementations
The distance metric
$\text{obs}(\mathbf{x},\mathbf{x}^{*})\;\coloneqq\;\sum_{\text{A}\in\text{Attr}}m\Big{(}x_{\text{A}},\,x^{*}_{\text{A}}\Big{)}/n$ (20)
is adopted from early work on counterfactual explanations (see e.g. Wachter et al. (2017)). We note that $x_{\text{A}}$ are standardized logits, so we do not require to scale the distances. The metric
$\text{plausible}(\mathbf{x}^{*})\;\coloneqq\;-\sum_{\text{A}\in\text{Attr}}\log p\left(x^{*}_{\text{A}}\;|\;\mathbf{x}^{*}_{\text{pa}(\text{A})}\right)/n$ (21)
measures the probability density of the attributes, which can be accessed for normalizing flows that we trained on the observed data set. Finally,
$\text{causal}(\mathbf{x},\mathbf{x}^{*})\;\coloneqq\;\sum_{\text{A}\in\text{Attr}}m\Big{(}f_{\text{A}}^{-1}(\mathbf{x}_{\text{pa}(\text{A})},x_{\text{A}}),\,f_{\text{A}}^{-1}(\mathbf{x}^{*}_{\text{pa}(\text{A})},\,x^{*}_{\text{A}})\Big{)}/n$ (22)
measures the distance of all latents for the structural model. We note that this loss corresponds to the distance term (the left summand) in 5, restricted to attribute variables. Since we do not have access to ground truth structural equations $(f_{1},f_{2},...,f_{n})$ in CelebA, we use the ones that were trained on the data set. We furthermore note that incorporating the antecedent variable into the loss is not an issue either, because it is fixed for all methods. For the deep non-causal explanation method, we only obtain the counterfactual image, without explicit access to the attribute variables. In order to extract those, we use the (standardized) logits of classifiers that were trained to predict the attributes from the image.
We obtain the numbers in table Tab. 1 as follows:
We sample a factual data point $\mathbf{x}=\mathbf{F}(\mathbf{u}),\mathbf{u}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Then, we sample an attribute uniformly, i.e.
@@ C.1 Metric Implementations
We then compute the counterfactual $\mathbf{x}^{*}$ to evaluate all three loss function 20, 21 and 22. This process is repeated $500$ times. The final reported scores are the arithmetic means over the individual metrics, including ±$\pm 1\text{std}$.
Appendix D Additional Experiments
D.1 Morpho-MNIST
Figure 11: We run Fig. 4 (b) multiple times, fixing $w_{I}=w_{\text{Img}}=1$ and changing only $w_{T}$. We see that the backtracking solution approaches the interventional solution (see Fig. 4 (b)) as we increase $w_{T}$, thus preserving the value of thickness more as we increase weight. We note that the left-most plot ($w_{T}=0$) corresponds to Fig. 4 (c).
D.2 CelebA

### v4

@@ A.2 Formal Derivation of DeepBC
We derive 7 and 5 from the three-step-procedure of backtracking counterfactuals (see § A.1) as follows:
@@
    We observe that 17 corresponds to performing stochastic DeepBC 6.
@@
    We conclude by noting that $\text{arg}\;\underset{\mathbf{u}^{\prime}}{\text{max}}\;\text{log}\;p(\mathbf{u}^{\prime}\;|\;\mathbf{F}^{-1}(\mathbf{x}),\mathbf{x}^{*}_{S})$, composed with $\mathbf{F}$, is equivalent to 7.
A.3 Derivation of 7
As a result of the linearization of $\mathbf{F}$, $\mathcal{L}(\mathbf{u}^{\prime};\mathbf{u},\mathbf{x}^{*}_{S})$ in 11 simplifies to
@@ B.1 Technical Details and Comments for the DeepBC Optimization Algorithm
In practice, we implement 14 as follows
@@
where $\dagger$ denotes Moore-Penrose pseudoinverse. We employ 19 rather than 14 for the reason of numerical stability. The main computational bottleneck in Alg. 1 is the computation of the pseudoinverse $(\lambda^{-1}\mathbf{W}+\mathbf{J}_{S}^{\top}\mathbf{J}_{S})^{\dagger}$ in 19, which comes at a cost of $\mathcal{O}(\#\text{it}\cdot\text{dim}(\mathbf{u})^{3})$, compared to $\mathcal{O}(\#\text{it}\cdot\text{dim}(\mathbf{u}))$ for gradient descent. We note, however, that the dimensionality of the latent space is typically not very large in our experiments. The maximum dimension is $516$ for CelebA, due to $4$ attributes and $512$-dimensional latent space of the VAE. We also stress that $\mathbf{J}_{S}$ is sparse (many 00 entries) when $S$ covers attribute variables, because the $512$-dimensional latent vector is not upstream of any attribute. We do not run experiments where many variables are upstream of the antecedent variable and stress that this may affect the performance of Alg. 1.
@@
Adam’s convergence highly depends on the choice of learning rate. We suspect that this is due to the poorly conditioned Hessian that comes from choosing large $\lambda$. However, large $\lambda$ is required in order to (at least approximately) fulfill the constraint in 7 (see top row in the table of Fig. 9).
Figure 9: The figures show the sum of penalty losses 11 over all points in Fig. 4 (b) for optimizing it for $200$ iterations on a $\text{log}_{10}$ scale. Comparison of the Adam optimizer with various learning rates in comparison to constraint linearization (Alg. 1) for different choices of penalty parameter $\lambda$. The table shows the sum of constraint errors $\left\lVert\mathbf{F}_{S}(\mathbf{u}^{\prime})-\mathbf{x}^{*}_{S}\right\rVert_{2}^{2}$ after $200$ iterations (How well the constraints are fulfilled).
@@ Appendix C Ground Truth Structural Equations in Morpho-MNIST
Appendix D Additional Experiments
D.1 Morpho-MNIST
Figure 11: (a) We run Fig. 4 (b) multiple times, fixing $w_{I}=w_{\text{Img}}=1$ and changing only $w_{T}$. We see that the backtracking solution approaches the interventional solution (see Fig. 4 (b)) as we increase $w_{T}$, thus preserving the value of thickness more as we increase weight. We note that the left-most plot ($w_{T}=0$) corresponds to Fig. 4 (c). (b) DeepBC can be extended to more general backtracking conditionals (corresponding distance functions plotted below each subplot) that lead to different solutions.
D.2 CelebA
D.2.1 Additional Qualitative Experiments
D.2.2 Quantitative Experiments
In addition to the qualitative experiments in § 4.2 and § D.2.1, we demonstrate quantitative experiments where we assess three metrics over a large sample of generated counterfactuals for the different methods. In addition to baselines 1 and 2 from the main text (§ 4.2), we assess a deep counterfactual explanation method 10. Specifically, we use an image regressor ($f_{\hat{Y}}$), together with an unconditional image auto-encoder ($f_{X}$) to generate counterfactual explanations, as sketched in Section § 3.2. We do so for each attribute ($\hat{Y}$) separately. This corresponds in style to how Jacob et al. (2022); Rodríguez et al. (2021) obtain counterfactual explanations for images. This means that $f_{X}$ does not correspond to the reduced form of a causal model, and so $u_{X}=f^{-1}_{X}(x)$ corresponds to an unstructured embedding space for image data. For our experiments, we apply DeepBC in this unstructured latent space (where neither the attributes nor their causal relationships are modeled) and use classifiers to extract the high-level attributes from the counterfactual image. We also use DeepBC with a wrong graph to assess the effect of model missspecification on the result. The wrong graph used is shown in Fig. 12 (c).
Quantitative Evaluation Metrics.
[…]
The plausible metric measures the probability density of the generated counterfactual attributes, under the data distribution. In this sense, we can think of counterfactuals with high probability density (or low -log density) as being more plausible, i.e., being closer to the data manifold (see Karimi et al. (2022)). The obs metric measures the distance among attributes between factual realization and counterfactual, thereby evaluating closeness in the observation space (regardless of whether the causal laws are preserved). This is the metric that has been optimized in the original work on counterfactual explanations (Wachter et al., 2017). The causal metric measures the difference in exogenous variables between realization and counterfactual, which can be interpreted as the degree of preservation of the causal mechanisms: If changes in latent variables exceed what is minimally necessary under the assumption of retained mechanisms, it suggests that the counterfactual does not effectively preserve the causal mechanisms. Since we do not have access to ground truth structural equations $(f_{1},f_{2},...,f_{n})$ in CelebA, we use the ones that were trained on the data set. We furthermore note that incorporating the antecedent variable into the loss is not an issue either, because it is fixed for all methods. For the deep non-causal explanation method, we only obtain the counterfactual image, without explicit access to the attribute variables. In order to extract those, we use the (standardized) logits of classifiers that were trained to predict the attributes from the image.
We obtain the numbers in table Tab. 1 as follows: We sample a factual data point $\mathbf{x}=\mathbf{F}(\mathbf{u}),\mathbf{u}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Then, we sample an attribute uniformly, i.e.
@@ D.2.2 Quantitative Experiments
We then compute the counterfactual $\mathbf{x}^{*}$ to evaluate all three loss functions 21. This process is repeated $500$ times. The final reported scores are the arithmetic means over the individual metrics, including ±$\pm 1\text{std}$.
Table 1: Quantitative Evaluations for CelebA. Three considered metrics $\pm$ standard deviation of the different baselines for $500$ iterations. Lower values are better, best per metric is highlighted in bold. CE stands for counterfactual explanation (see § 2.4). The two distance functions assessed ($m$) are SQU for squared distance, and ABS for absolute distance.
Metric m Tabular CE Interventional Deep CE Wrong graph DeepBC
Results. We quantitatively assess multiple properties of sparse mode DeepBC in comparison to the baselines in Tab. 1. While the deep counterfactual explanation method 10 achieves the best result on the plausible score, the interventional approach preserves the attribute values the best (obs score). This is likely due to the fact that counterfactuals on leaf nodes of the causal graph in Fig. 6 do not change the values of their parents at all, in contrast to the other methods. Finally, as expected, DeepBC obtains the best score on the causal metric, independent of distance function $m$. This can be explained due to the fact that DeepBC always respects downstream effects (in comparison to other methods), as exemplified in Fig. 6.